OpenStudy (hotpinkgirl):

Ball A and ball B are rolled so that they will collide. does the data in the table below demonstrate the law of conservation of momentum? A. yes, because the total momentums before and after the collision are equal. B. yes, because the total momentums before and after the collision are unequal C. no, because the total momentums before and after the collision are equal D. no, because the total momentums before and after the collision are unequal

OpenStudy (hotpinkgirl):



OpenStudy (hotpinkgirl):

the last one says: total momentom of both balls after collision

OpenStudy (hotpinkgirl):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm not 100% familiar with physics, but I would think that ball A would be +3 kg m/s and ball B would be -2 kg m/s. The negative is to mean ball B is going in the opposite direction. So adding them gives +3 + (-2) = +1 which means that the two masses will stick together and have momentum of 1 kg m/s

jimthompson5910 (jim_thompson5910):

then again I'm not an expert in physics, so I'm not sure

OpenStudy (hotpinkgirl):

so it would be D????

OpenStudy (hotpinkgirl):

@jim_thompson5910 so it would be D????

jimthompson5910 (jim_thompson5910):

yeah something seems off, so I'm thinking D too. But I'd get a second opinion. Preferably someone who knows their physics.

OpenStudy (hotpinkgirl):

@shrutipande9

OpenStudy (shrutipande9):

momentum after collision of both balls is 6 each?

OpenStudy (hotpinkgirl):

so its A right @shrutipande9

OpenStudy (shrutipande9):

@Abhisar

OpenStudy (shrutipande9):

r u sure its equal?

OpenStudy (abhisar):

Law of conservation of linear momentum says that the total momentum of the system before the collision is equal to the linear momentum of the system after collision. Now, from the table, we see that the total momentum of the system before the collision is 3+2 = 5 Kgm/s while total linear momentum after the collision is 6 Kgm/s So, what do you think should be the correct option?

OpenStudy (hotpinkgirl):

it is D right @Abhisar

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (hotpinkgirl):

Thank you soo much!!!!!!!!!!

OpenStudy (abhisar):

This question is unrelated to Biology section so use Physics section for similar questions.

OpenStudy (hotpinkgirl):

ok

OpenStudy (abhisar):

Welcome (-: \(\color{olive}{\Huge\widehat{\underset{^\backslash \textrm{____} ^/}{\left\{\underset{\color{#aa1111}{\leftrightsquigarrow}}{\!\!\!\color{red}{\bigodot}\,_{\Large\color{#0f500f}{\bullet\bullet}}\Huge\color{red}{\bigodot}\!\!\!}\right\}}}}\)